Workflow Team Meeting - March 31 4PM CERN, 10 FNAL time

Vidyo Link


  • FNAL: Jen, Gaston, Jesus, Scarlet, SeangChan, Jorge
  • US: Matteo, Alli
  • CERN : Paola, Dima, JeanRoc, Welcome Sebastian - new transfer person


  • Jorge to Columbia April 15-May 2, Talk on April 27
  • SeangChan taking 1/2 day next Thurs and Fri off next week

News - Dima

  • The next reprocessing big tests are getting going, big test workflow just injected 150 priority so it should take over everything, so we need to pay attention to this workflow we should be running ~70K/4 cores in parallel if this goes well we will know that we are ready to run multicore for real
    • keep an eye on pdmvserv_task_TOP-RunIISpring16DR80-00001__v1_T_160331_151408_3872

  • we need to prioritize running at the T0, we need to fully saturate the T0, 13K jobs, and jobs were crashing,
    • we have a problem, the T0 was down there was something wrong with the production status metric
    • vlimant_BTV-RunIISummer15GS-Backfill-00048_00275_v0__160330_122825_8568 - is the test workflow
  • April fools reprocessing will hopefully start on monday, we still need to prestage data, it's not ready yet, Dima is working on cleaning things up by the end of the day so things can be prestaged. We will be using high priority transfers to get things staged quickly.

3 top issues affecting production

  • ReReco Workflow not being picked up by Unifed
  • just testing and waiting
  • stuck workflows - there was a problem with T0_CH_CERN and Alan helped us clear them out

Site support - Gaston

2016-03-25 00:00:01 T2_EE_Estonia x      
2016-03-25 00:00:01 T2_IN_TIFR x      
2016-03-30 00:00:01 T2_PK_NCP     x  
2016-03-30 00:00:01 T2_UK_London_Brunel x      

  • problem with the production script, all the T3's are listed as on for production
  • there is a change in how the production metric will be used, Gaston will work with the team to get it set properlly

Transfers - Jorge

  • No new missing files at KIT so thank you Alan and Jorge! Old workflows may still give us missing files so the workflow sandboxes will be effected, March 16 is the day the changes went in, so things older than that we may have to redo, things newer than that should be fine


  • All work backlog is current to the past month!





  • NA


  • fabozzi_Commissioning2015-Cosmics-Boff-01Mar2016_763p2_160302_100618_8061 needs to be cloned this one got lost in file invalidation land
    • it has been cloned: jen_a_Commissioning2015-Cosmics-Boff-01Mar2016_763p2_160331_172001_8114

Store Results


Agent Issues

Agent redeployment

RequestMgr2 Migration

Merging Scripts

RelVal Andrew


-- JenniferAdelmanMcCarthy - 2016-03-29

Edit | Attach | Watch | Print version | History: r6 < r5 < r4 < r3 < r2 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r6 - 2016-03-31 - JenniferAdelmanMcCarthy
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    CMSPublic All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright & 2008-2021 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback